Question

Using the information from above, with p′=0.143, q′=0.857, and n=350, what is the 95% confidence interval for the proportion of the population who play games on their phones?

z0.10 z0.05 z0.025 z0.01 z0.005
1.282 1.645 1.960 2.326 2.576

Use the table of common z-scores above.

  • Round the final answer to three decimal places.

Provide your answer below:
Answer #1

Solution :

Given that,

n = 350

Point estimate = sample proportion = \hat p = 0.143

1 - \hat p = 0.857

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.960 * (\sqrt((0.143*0.857) / 350)

= 0.037

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.143 - 0.037 < p < 0.143 + 0.037

0.106 < p < 0.180

The 95% confidence interval for the population proportion p is : ( 0.106 , 0.180 )